OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (aravindg):

well for first one use the property that (a+b)(a-b)=a^2-b^2

OpenStudy (aravindg):

for second one use (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

k thanks

OpenStudy (aravindg):

can you click best response ?

OpenStudy (aravindg):

@Anna07